OpenStudy (anonymous):

Find the integral ( from 0 to 2): qdq / (8q +5) im trying to use u substitution.

OpenStudy (anonymous):

\[\int_0^2\frac{q}{8q+5}~dq\] Let \(u=8q+5\), so that \(\dfrac{1}{8}du=dq\). Note that \(u=8q+5~\iff~q=\dfrac{u-5}{8}\). Now substitute into the integral, keeping in mind the change in limits: \[\frac{1}{8}\int_5^{21}\frac{\frac{u-5}{8}}{u}~du\\ \frac{1}{64}\int_5^{21}\frac{u-5}{u}~du\\ \frac{1}{64}\int_5^{21}\left(1-\frac{5}{u}\right)~du\\ \]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're welcome